Determine the resultant of the three forces acting on the gusset plate of a truss.


1
Expert's answer
2021-06-07T08:31:57-0400


Resolving the forces into horizontal and vertical components:

"∑V(↑+) = 6200-4000sin45-3800sin50 = 460.604 \;N \\

∑H→+=4000cos45-3800cos50=385.834 \;N"

The resultant force is

"R=460.6042+385.8342 =600.85 \;N"